OpenStudy (anonymous):

What is the area of a hexagon with the radius of 5 inches?

OpenStudy (jdoe0001):

how many sides does a hexagon have anyway?

OpenStudy (anonymous):

6 I think?

OpenStudy (jdoe0001):

yeap, it has 6, HEXA = 6, thus \(\textit{area of a regular polygon}=\cfrac{1}{4}ns^2\ cot\left( \frac{180}{n} \right)\qquad \begin{cases} n=\textit{number of sides}\\ s=\textit{length of one side} \end{cases}\)

OpenStudy (anonymous):

Would it be 64.95?

OpenStudy (jdoe0001):

what did you get for the cotangent anyway?

OpenStudy (anonymous):

The what?

OpenStudy (jdoe0001):

http://www.mathopenref.com/polygonregulararea.html

OpenStudy (jdoe0001):

\(\textit{area of a regular polygon}=\cfrac{1}{4}ns^2\ cot\left( \frac{180}{n} \right)\qquad \begin{cases} n=\textit{number of sides}\\ s=\textit{length of one side} \end{cases}\) plug in those values, see what you get